# Marginal phase transition in decorated single-chain Ising models

### v1

@@ Abstract
Since Ernst Ising’s proof one century ago, it has been well-known that phase transition at finite temperature does not exist in the Ising model with short-range interactions in one dimension. Yet, little is known about whether this forbidden transition could be approached arbitrarily closely—at fixed finite temperature. To explore such asymptoticity, the notion of marginal phase transition (MPT) was introduced recently and spontaneous MPT was successfully found in decorated ladder Ising models. On the other hand, in the presence of a magnetic field, narrow phase crossover termed as pseudo-transition was found in decorated single-chain Ising models with strong geometric frustration; it is thus imperative to know whether the pseudo-transition could be transformed to approach a genuine transition at fixed finite temperature $T_{0}$ arbitrarily closely, i.e., being the MPT. Here, I reveal the existence of the field-induced MPT in decorated Ising chains, in which $T_{0}$ is determined by the interactions involving only the decorated parts and the magnetic field, while the crossover width $2\delta T$ is independently, exponentially reduced ($\delta T=0$ means a genuine transition) by the previously neglected ferromagnetic interaction between the ordinary spins on the chain backbone. Furthermore, I show that the MPT can be realized even in the decorated Ising chains without geometric frustration because the magnetic field itself can induce previously unnoticed hidden spin frustration. These findings manifest that MPT is essentially the buildup of coherence in preformed crossover of any local states, making the doors wide open to the engineering and utilization of MPT as a new paradigm for exploring exotic phenomena and 1D device applications.
The Ising model Ising1925 describes collective behaviors such as phase transitions and critical phenomena in various physical, biological, economical, and social systems Mattis_book_08_SMMS ; Mattis_book_1985 ; Baxter_book_Ising ; 003_Huang_08_book . The one-dimensional (1D) Ising model on a decorated single chain is generally defined as $H=H_{\mathrm{ordinary}}+\sum_{i}H^{(i)}_{\mathrm{decorated}}$, where
@@
describes the ordinary single chain without the decoration (Fig. 1a) with $\sigma_{i}=\pm 1$ standing for the $i$th ordinary spin—in fact, it can be used to describe any two-value system, eg. open or close in neural networks Schneidman2006 , yes or no in voting. $N$ is the total number of the ordinary spins and $\sigma_{N+1}\equiv\sigma_{1}$ (i.e., the periodic boundary condition). $J$ is the interaction between nearest-neighboring ordinary spins. $h$ depicts the magnetic field and $\mu_{a}$ the magnetic moment of the ordinary spins. $H^{(i)}_{\mathrm{deorated}}$ describes the decorated part in between the $i$th and $(i+1)$th ordinary spins (Fig. 1b), which can be any finite-size subsystem as long as it couples to the two nearest ordinary spins by the Ising-type interactions only. To date, the simplest $H^{(i)}_{\mathrm{deorated}}$ considered in the literature of “pseudo-transition” 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter is the Ising diamond (Fig. 1d) given by 016_Strecka_book_chapter
@@
To search for MPT, which resembles a genuine first-order transition with large latent heat at fixed finite temperature $T_{0}$, in the decorated single-chain Ising models, it is instrumental to start with a comparative review of (i) the previous studies of the decorated Ising chains with pseudo-transition in the presence of magnetic field 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter and (ii) the recent investigations of the decorated Ising ladders with MPT in the absence of the magnetic field Yin_MPT ; Yin_icecreamcone ; Hutak_PLA_21_trimer . In particular, we ask questions as to what the pseudo-transition research has done and has not done, compared with what we have learned from the spontaneous MPT in the ladder. This task is greatly simplified by a recent summary 017_Krokhmalskii_PA_21_3-previous-chains_effective_model of the pseudo-transition research in the effective Hamiltonian approach, where tracing out the decorated parts results in the ordinary Ising-chain model with temperature-dependent parameters $J_{\mathrm{eff}}(T)$ and $h_{\mathrm{eff}}(T)$ in place of $J$ and $h$ in Eq. (1). Two key conclusions about the existence of the pseudo-transition were reached 017_Krokhmalskii_PA_21_3-previous-chains_effective_model : (1) $h_{\mathrm{eff}}(T)$ must experience the sign change as a function of temperature $T$ and $T_{0}$ is determined by $h_{\mathrm{eff}}(T)=0$. (2) The decoration was done to create geometrical frustration so that the system’s first low-lying excited state has much higher degeneracy and just slightly higher energy than the ground state, then an entropy-driven crossover between them would occur at finite temperature 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter . This physics of phase crossover is rather generic Miyashita_10_review_frustration ; what is challenging is how to make the crossover ultra-narrow. It was found that $T_{0}\to 0$ as the pseudo-transition is “tracked down in the critical point of the standard Ising-chain model at $h=0$ and $T=0$” 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; therefore, the pseudo-transition does not appear to support MPT at fixed finite $T_{0}$. The fact that spontaneous pseudo-transition in the decorated single-chain Ising models is prohibited by symmetry (namely the Ising model is invariant to the flipping of all the spins for $h=0$) motivated the discovery of the decorated Ising ladders as the paradigm for spontaneous MPT Yin_MPT ; Yin_icecreamcone . Now that the knowledge of MPT has become available, we understand that the following key pieces of information were missing in the previous studies of pseudo-transition and we are able to quickly find the solution:
@@
1) On the description of the targeted phenomenon.—The crossover width $2\delta T$ was never clearly defined and expressed by the model parameters (note that $\delta T=0$ means a genuine phase transition) for the pseudo-transition, while it was always presented in terms of specific heat, entropy, magnetic susceptibility, or overall magnetization 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter ; these physical quantities are derivatives of the free energy with respect to the global parameter $T$ or $h$ and thus depend on the details of the model. For spontaneous MPT in decorated Ising ladders, the on-rung parent-spin correlation function was identified as the order parameter (OP) that has a well-defined value space of $[-1,1]$ with the value 00 meaning $T_{0}$ and its inverse slope at $T_{0}$ meaning $\delta T$, characterizing the MPT as an abrupt change in the OP between nearly $-1$ to nearly $+1$ Yin_MPT ; Yin_icecreamcone . Moreover, the general form of the OP does not depend on the details of the model; its mathematical derivation and numerical computation can be easily carried out. Such an OP provides an accurate, convenient, and microscopic description of the MPT; its identification greatly accelerated the search for MPT. Here for the decorated Ising chains, the OP that has the same features is $\langle\sigma_{i}\rangle$, the magnetization of the ordinary spins on the chain backbone (not the overall magnetization that includes the decorated parts): Its sign change and zero value at $T_{0}$ is consistent with the behavior of $h_{\mathrm{eff}}(T)$; its inverse derivative at $T_{0}$ defines $\delta T=\left(\partial\langle\sigma_{i}\rangle/\partial T\right)^{-1}_{T=T_{0}}$ (Fig. 5a).
@@
2) On the model Hamiltonian.—Surprisingly, the $J$ term—the Ising interaction between the ordinary spins on the chain backbone (red bonds in Figs. 1a, 1b, 1e, 1f)—was neglected in the previous studies of pseudo-transition 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter , that is, Fig. 1c was studied instead of the more general Fig. 1b. A possible reason for the omission of $J$ could be that the standard geometric frustration from the triangles formed by the antiferromagnetic bonds is more obvious for $J=0$, as shown in Fig. 1d for the Ising diamond chain, the hitherto simplest model with pseudo-transition. However, the MPT for the Ising ladders tells us that the on-leg decoration (which controls $\delta T$) can be done independently of the on-rung decoration (which controls $T_{0}$in_MPT ; Yin_icecreamcone . We shall use the $J\neq 0$ Ising diamond chain model (Fig. 1e) to show that similarly, the $J$ term independently exponentially reduces the crossover width $2\delta T$ for fixed finite $T_{0}$, since $J$ has no effect on $h_{\mathrm{eff}}(T)$ but is a separate addend in $J_{\mathrm{eff}}(T)$, i.e., $J_{\mathrm{eff}}(T,J)=J+J_{\mathrm{eff}}(T,0)$ (see Eq. (16) in the Method).
@@
3) On the underlying mechanism.—Only the frustration of geometric frustration type (namely the lattice has frustration regardless of the magnetic field) was considered for the pseudo-transition. Using the effective Hamiltonian approach to study the same spontaneous MPT in the decorated Ising ladders with geometrically frustrated triangles as Refs. Yin_MPT ; Yin_icecreamcone , Hutak et al. conjectured that geometric frustration is not necessarily a prerequisite for the realization of spontaneous MPT in such ladders Hutak_PLA_21_trimer . Although this conjecture has not been proved for the ladder yet, we were inspired to extend the conjecture to the decorated single-chain Ising models, because it was recently emphasized Yin_g that the magnetic field on its own could induce hidden spin frustration in ferrimagnet-like systems without geometric frustration Bell_JPC_74_Ising_ferri_1 . It suffices to prove the conjecture with one example of the systems without geometric frustration; here we show that it is true in the Ising diamond chain for $J_{2}=0$ (Fig. 1f), where the triangles formed by the $J$ and $J_{1}$ bonds are not frustrated because $J>0$ is ferromagnetic. The hidden high degeneracy generated by the magnetic field, not by geometric frustration, is explained in the ground-state phase diagram of the Ising diamond chain (the red circle on the red dashed line in Fig. 2).
@@
These findings thoroughly expose the mathematical structure of MPT in the 1D Ising models. One can use various decorations to yield even broad phase crossovers and then use interactions that enhance the coherence of the order parameter to turn the broad phase crossover to be the MPT, reminiscent of the notion of preformed pairs and their coherence buildup in the field of high-temperature superconductivity Emery_Nature_95 . Given the physical effects that the MPT resembles a genuine first-order phase transition with large latent heat and the fact that the Ising model has already been implemented in electronic circuits Ising_FPGA and optical networks Pierangeli_IsingMachine_PRL19 , the MPT-based 1D devices for thermal applications appear to be feasible. The features that $T_{0}$ and $2\delta T$ can be independently controlled by different parameters and different decoration methods could be attractive in engineering 1D thermal sensors, for example. The doors to the engineering and utilization of MPT are now wide open.

### v2

# Paradigm for approaching the forbidden phase transition in the one-dimensional Ising model at fixed finite temperature: Single chain in a magnetic field

@@ Abstract
In a previous paper [Weiguo Yin, Phys. Rev. Res. 6, 013331 (2024)], the forbidden spontaneous phase transition in the one-dimensional Ising model was found to be approachable arbitrarily closely in decorated ladders by ultra-narrow phase crossover (UNPC) at a given finite temperature $T_{0}$ with the crossover width $2\delta T$ reduced exponentially, which resemble a genuine first-order transition with large latent heat. Here, I reveal that the forbidden phase transition can be approached at fixed $T_{0}$ as well in decorated single-chain Ising models in the presence of a magnetic field, in which $T_{0}$ is determined by the interactions involving only the decorated parts and the magnetic field, while $2\delta T$ is independently, exponentially reduced ($\delta T=0$ means a genuine transition) by restoring the ferromagnetic interaction between the ordinary spins on the chain backbone—which was neglected in the previous studies of pseudo-transition—thus manifesting that this asymptoticity to the forbidden transition is essentially the buildup of coherence in preformed crossover of local states. Furthermore, I show that the UNPC can be realized even in the absence of the conventional geometric frustration because the magnetic field itself can induce previously unnoticed hidden spin frustration. These findings make the doors wide open to the engineering and utilization of UNPC as a new paradigm for exploring exotic phenomena and 1D device applications.
I Introduction
The textbook Ising model describes collective behaviors such as phase transitions and critical phenomena in various physical, biological, economical, and social systems Mattis_book_08_SMMS ; Mattis_book_1985 ; Baxter_book_Ising ; 003_Huang_08_book . Since Ernst Ising’s proof one century ago Ising1925 , it has been well-known that phase transition at finite temperature does not exist in the Ising model with short-range interactions in one dimension Cuesta_1D_PT . Yet, little is known about whether this forbidden transition could be approached arbitrarily closely—at fixed finite temperature $T_{0}$—until recently such asymptoticity was successfully found in decorated ladder Ising models in the absence of a magnetic field Yin_MPT ; Yin_icecreamcone . On the other hand, in the presence of a magnetic field, ultra-narrow phase crossover termed as “pseudo-transition” was found in decorated single-chain Ising models with strong geometric frustration 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter ; however, there is no hint to the question of how to make the crossover width $2\delta T$ narrower and narrower while keeping $T_{0}$ unchanged, since $2\delta T$ was not even defined in terms of the model parameters. The independent control of $T_{0}$ and $2\delta T$ by different interactions will not only help push the limit in our understanding of phase transitions arbitrarily close to the forbidden regime, but also provide promising potentials in technology applications. It is thus imperative to explore whether and how the forbidden in-field phase transition could be approached arbitrarily closely at a given $T_{0}$ in the single-chain Ising models.
The one-dimensional (1D) Ising model on a decorated single chain is generally defined as $H=H_{\mathrm{ordinary}}+\sum_{i}H^{(i)}_{\mathrm{decorated}}$, where
@@ I Introduction
describes the ordinary single chain without the decoration (Fig. 1a) with $\sigma_{i}=\pm 1$ standing for the $i$th ordinary spin—in fact, it can be used to describe any two-value system, eg. open or close in neural networks Schneidman2006 , yes or no in voting. $N$ is the total number of the ordinary spins and $\sigma_{N+1}\equiv\sigma_{1}$ (i.e., the periodic boundary condition). $J$ is the interaction between nearest-neighboring ordinary spins. $h$ depicts the magnetic field and $\mu_{a}$ the magnetic moment of the ordinary spins. $H^{(i)}_{\mathrm{deorated}}$ describes the decorated part in between the $i$th and $(i+1)$th ordinary spins (Fig. 1b), which can be any finite-size subsystem as long as it couples to the two nearest ordinary spins by the Ising-type interactions only. To date, the simplest $H^{(i)}_{\mathrm{deorated}}$ considered in the literature of “pseudo-transition” is the Ising diamond 016_Strecka_book_chapter (Fig. 1d) given by
@@
It is instrumental to start with a comparative review of (i) the previous studies of the decorated Ising chains with pseudo-transition in the presence of magnetic field 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter and (ii) the recent investigations of the decorated Ising ladders in the absence of the magnetic field Yin_MPT ; Yin_icecreamcone ; Hutak_PLA_21_trimer . In particular, we ask questions as to what the pseudo-transition research has done and has not done, compared with what we have learned from the spontaneous ultra-narrow phase crossover in the ladder. This task is greatly simplified by a recent summary 017_Krokhmalskii_PA_21_3-previous-chains_effective_model of the pseudo-transition research in the effective Hamiltonian approach, where tracing out the decorated parts results in the ordinary Ising-chain model with temperature-dependent parameters $J_{\mathrm{eff}}(T)$ and $h_{\mathrm{eff}}(T)$ in place of $J$ and $h$ in Eq. (1). Two key conclusions about the existence of the pseudo-transition were reached 017_Krokhmalskii_PA_21_3-previous-chains_effective_model : (1) $h_{\mathrm{eff}}(T)$ must experience the sign change as a function of temperature $T$ and $T_{0}$ is determined by $h_{\mathrm{eff}}(T)=0$. (2) The decoration was done to create geometrical frustration so that the system’s first low-lying excited state has much higher degeneracy and just slightly higher energy than the ground state, then an entropy-driven crossover between them would occur at finite temperature 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter . This physics of phase crossover is rather generic Miyashita_10_review_frustration . To make the crossover ultra-narrow, it is of normal practice to place the system close to the critical point or the phase boundary in the zero-temperature phase diagram. It was found that $T_{0}\to 0$ as the pseudo-transition is “tracked down in the critical point of the standard Ising-chain model at $h=0$ and $T=0$” 017_Krokhmalskii_PA_21_3-previous-chains_effective_model . In other words,
[…]
in this traditional paradigm of realizing the ultra-narrow phase crossover by approaching the zero-temperature phase boundary of two competing phases. Therefore, the pseudo-transition does not appear to support our goal of approaching the forbidden phase transition at fixed finite $T_{0}$. Now that the knowledge of how to realizing the goal in the Ising ladders has become available Yin_MPT ; Yin_icecreamcone , we understand that the following key pieces of information were missing in the previous studies of pseudo-transition and we are able to quickly find the solution in the term of “coherent pseudo-transition” (CPT) that unifies the ultra-narrow phase crossovers in both the ladder and the single-chain Ising models:
@@
1) On the description of the targeted phenomenon.—The crossover width $2\delta T$ was never clearly defined and expressed by the model parameters (note that $\delta T=0$ means a genuine phase transition) for the pseudo-transition, while it was always presented in terms of specific heat, entropy, magnetic susceptibility, or overall magnetization 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter ; these physical quantities are derivatives of the free energy with respect to the global parameter $T$ or $h$ and thus depend on the details of the model. For spontaneous CPT in decorated Ising ladders, the on-rung parent-spin correlation function was identified as the order parameter (OP) that has a well-defined value space of $[-1,1]$ with the value 00 meaning $T_{0}$ and its inverse slope at $T_{0}$ meaning $\delta T$, characterizing the CPT as an abrupt change in the OP between nearly $-1$ to nearly $+1$ Yin_MPT ; Yin_icecreamcone . Moreover, the general form of the OP does not depend on the details of the model; its mathematical derivation and numerical computation can be easily carried out. Such an OP provides an accurate, convenient, and microscopic description of the CPT; its identification greatly accelerated the search for CPT. Here for the decorated single-chain Ising models, the OP that has the same features is $\langle\sigma_{i}\rangle$, the magnetization of the ordinary spins on the chain backbone (not the overall magnetization that includes the decorated parts): Its sign change and zero value at $T_{0}$ is consistent with the behavior of $h_{\mathrm{eff}}(T)$; its inverse derivative at $T_{0}$ defines $\delta T=\left|\partial\langle\sigma_{i}\rangle/\partial T\right|^{-1}_{T=T_{0}}$ (Fig. 5a).
@@
2) On the model Hamiltonian.—Surprisingly, the $J$ term—the Ising interaction between the ordinary spins on the chain backbone (red bonds in Figs. 1a, 1b, 1e, 1f)—was neglected in the previous studies of pseudo-transition 005_Galisova_PRE_15_double-tetrahedral-chain ; 007_Torrico_PRA_16_Ising-XYZ-diamond-chain ; 009_review_Souza_SSC_18_Ising-XYZ-diamond-chain_double-tetrahedral-chain-spin-electron ; 010_Carvalho_JMMM_18_Ising-XYZ-diamond-chain_quantum-entanglement ; 011_Rojas_BJP_20_Ising-Heisenberg-tetrahedral_diamond ; 013_Rojas_PRE_19_previous_4_models ; 014_Rojas_JPC_20_Ising–Heisenberg_spin-1-double-tetrahedral-chain ; 015_Strecka_APPA_20_Ising-diamond-chain ; 015-7_Canova_CzechoslovakJP_04_Ising–Heisenberg_diamond_chain ; 015-8_Canova_JPC_06_Ising–Heisenberg-spin-S-diamond-chain ; 017_Krokhmalskii_PA_21_3-previous-chains_effective_model ; 016_Strecka_book_chapter , that is, Fig. 1c was studied instead of the more general Fig. 1b. A possible reason for the omission of $J$ could be that the standard geometric frustration from the triangles formed by the antiferromagnetic bonds is more obvious for $J=0$, as shown in Fig. 1d for the Ising diamond chain, the hitherto simplest model with pseudo-transition. However, the CPT for the Ising ladders tells us that the on-leg decoration (which controls $\delta T$) can be done independently of the on-rung decoration (which controls $T_{0}$in_MPT ; Yin_icecreamcone . We shall use the $J\neq 0$ Ising diamond chain model (Fig. 1e) to show that similarly, the $J$ term independently exponentially reduces the crossover width $2\delta T$ for fixed finite $T_{0}$, since $J$ has no effect on $h_{\mathrm{eff}}(T)$ but is a separate addend in $J_{\mathrm{eff}}(T)$, i.e., $J_{\mathrm{eff}}(T,J)=J+J_{\mathrm{eff}}(T,0)$ (see Eq. (17) in the Method).
@@
3) On the underlying mechanism.—Only the frustration of geometric frustration type (namely the lattice has frustration regardless of the presence or absence of the magnetic field) was considered for the pseudo-transition. It was recently emphasized Yin_g that the magnetic field on its own could induce hidden spin frustration in ferrimagnet-like systems without geometric frustration Bell_JPC_74_Ising_ferri_1 . Here we show the existence of CPT in the Ising diamond chain for $J_{2}=0$ (Fig. 1f), where the triangles formed by the $J$ and $J_{1}$ bonds are not frustrated because $J>0$ is ferromagnetic. The hidden high degeneracy generated by the magnetic field, not by geometric frustration, is explained in the ground-state phase diagram of the Ising diamond chain (the red circle on the red dashed line in Fig. 2). An immediate improvement is the dramatic increase in $T_{0}$, e.g., by 2200% when $J_{2}/J_{1}$ is moved away from $1.95$ to 00 (note that $J_{2}/J_{1}=2$ sets the phase boundary), as shown in Fig. 3(b) and Figs. 4(b)(d).
@@
These findings thoroughly expose the mathematical structure of CPT as a generic mechanism for generating ultra-narrow phase crossover at fixed $T_{0}$ in the 1D Ising models. One can use various decorations to yield even very broad phase crossovers and then use interactions that enhance the coherence of the order parameter to turn the broad phase crossover to be the CPT—that is how CPT (coherent pseudo-transition) got its name—reminiscent of the notion of preformed pairs and their coherence buildup in the field of high-temperature superconductivity Emery_Nature_95 . Given the physical effects that the CPT resembles a genuine first-order phase transition with large latent heat and the fact that the Ising model has already been implemented in electronic circuits Ising_FPGA , optical networks Pierangeli_IsingMachine_PRL19 , and optical lattices Bernien_Nature_17_Rydberg , the CPT-based 1D devices for thermal applications appear to be feasible. The features that $T_{0}$ and $2\delta T$ can be independently controlled by different parameters and different decoration methods could be attractive in engineering 1D thermal sensors, for example. The doors to the engineering and utilization of CPT are now wide open.
II RESULTS AND DISCUSSIONS